# Learn while Unlearn: An Iterative Unlearning Framework for Generative Language Models

### v3

@@ Abstract.
Machine Unlearning, Natural Language Processing, Generative Language Model
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: ; ; booktitle: Preprintisbn: 978-1-4503-XXXX-X/2018/06
1. Introduction
@@
The rapid advancement of deep learning, particularly in Natural Language Processing (NLP), has led to the widespread deployment of powerful generative language models (GLMs) such as GPT-4 (Achiam et al., 2023), Claude 3 (Anthropic, 2024), and Google Gemini (Team et al., 2023). These models exhibit remarkable capabilities in text generation, summarization, and conversational AI, transforming numerous real-world applications. However, this progress has also raised critical concerns regarding data privacy, security, and compliance with evolving regulations. One of the most pressing challenges is the risk of unintentional memorization and leakage of sensitive information, such as personally identifiable data, proprietary content, and confidential records contained in training datasets (Carlini et al., 2021; Brown et al., 2022; Carlini et al., 2022).
@@
To mitigate these risks, regulatory frameworks like the General Data Protection Regulation (GDPR) (Voigt and Von dem Bussche, 2017) have introduced principles such as the “Right To Be Forgotten” (RTBF) (Mantelero, 2013; Villaronga et al., 2018), which necessitate the ability to remove or selectively forget specific information from trained machine learning models. This has driven extensive research into machine unlearning, a field that aims to ensure models can eliminate particular knowledge while maintaining overall functionality and performance.
@@
Figure 1. Difference among the generated sequences of (a) original model, (b) model unlearned by KUMPR (Jang et al., 2023) and (c) model unlearned by our method.
[…]
Comparing text sequences generated by three models in side-by-side panels. Panel (a) shows outputs from the original model, which indicates its knowledge of “Harry Potter”, panel (b) demonstrates outputs after unlearning via KUMPR, showing nonsensical repetition, and panel (c) displays outputs from the proposed ICU. Compared to (a) and (b), our proposed method (c) appears to reduce repetition and improve specificity, showing concise, contextually appropriate responses without redundant or irrelevant content.
@@
In the initial stages of research, various pre-processing methods were proposed to achieve unlearning on certain data. For instance, Kandpal et al. (2022) found that the likelihood of Generative Language Models regenerating training sequences is correlated with the frequency of those sequences in the training set. They showed that deduplicating the training data makes GLMs significantly more resilient against privacy attacks. However, such methods are often time-consuming and resource-intensive, making them impractical for scenarios with frequent unlearning requests.
@@
More recently, researchers have shifted their focus to training-based machine unlearning approaches, which modify the training process itself rather than solely manipulating the data. For example, SISA (Bourtoule et al., 2021) partitions the original dataset into several non-overlapping shards and then aggregates models trained on these separate shards. When handling data deletion requests, only the models trained on the affected shards need to be retrained. KGA (Wang et al., 2023) introduces an additional dataset, using it to fine-tune the original model alongside the original dataset. Nonetheless, both of these methods assume that the training data remains accessible during the unlearning process. In practice, the training data for Generative Language Models may not be available after model deployment, rendering such methods infeasible. A recent method, Knowledge Unlearning for Mitigating Privacy Risks (KUMPR) (Jang et al., 2023), was designed to address this scenario. [^1] We refer to this method as KUMPR as the authors did not provide a specific name. KUMPR reverses the training objective by maximizing the negative log-likelihood for target tokens and uses metrics to determine whether the model has “forgotten” a target sequence.
@@
  • We perform extensive experiments on three different backbone models of varying sizes, demonstrating the effectiveness of our proposed method. Our code is available at
    https://github.com/himalalps/ICU.
2. Related work
@@ 2.1. Machine unlearning
Machine unlearning, introduced by Cao and Yang (2015), aims to protect machine learning models from extraction attacks by removing specific data in such a way that the model behaves as if the data was never part of the training set. Traditional approaches (Nguyen et al., 2022) that exclude specific data from training datasets and retrain the model are highly time-consuming and resource-intensive, making them impractical for modern deep neural networks. Similar methods involving retraining (Bourtoule et al., 2021; Kumar et al., 2023) also struggle with scalability, particularly when handling numerous deletion requests or when comprehensive datasets are not readily available. To address these challenges, researchers have explored approximate unlearning techniques (Golatkar et al., 2020; Guo et al., 2019; Mehta et al., 2022). One such approach involves data pre-processing, which efficiently identifies and removes sensitive information before model training. Kandpal et al. (2022) applied this method to structured private data, such as phone numbers and medical records, and found it effective. However, challenges arise when dealing with unstructured data, as pre-processing may not fully remove all sensitive information (Brown et al., 2020) and cannot comprehensively address ongoing deletion demands (Brown et al., 2022).
@@
Recent studies (Jang et al., 2023; Kassem et al., 2023; Qu et al., 2024; Yao et al., 2023; Eldan and Russinovich, 2023) have focused on fine-tuning Generative Language Models to tackle machine unlearning challenges. Jang et al. (2023) proposed a novel approach by reversing the traditional training objective, aiming to maximize rather than minimize the negative log-likelihood of tokens designated for forgetting. Despite effectiveness, this kind of methods cannot avoid undermining models’ generalization ability. Other recent methods (Chen and Yang, 2023; Gao et al., 2024; Gu et al., 2024; Kassem et al., 2023; Wang et al., 2023; Maini et al., 2024; Yao et al., 2023; Li et al., 2024) use diverse techniques such as knowledge gap alignment and reinforcement learning. Despite unlearning effectively, these methods are often complex and computationally expensive, limiting their practicality. For instance, Gu et al. (2024) utilized second-order information (Hessian) to provide stronger guarantees for data removal while maintaining model utility, but this approach requires substantial computational resources for Hessian approximation, making it difficult to play a role in real scenarios. Pawelczyk et al. (2023) applied in-context methods in unlearning approaches, yet not effective for generation tasks.
2.2. Generative language models
@@
Generative Language Models are designed to understand, generate, and predict human language (Zhao et al., 2023), which have gained considerable attention in recent years (Jang et al., 2023; Kassem et al., 2023; Qu et al., 2024; Yao et al., 2023).
@@
Traditional language models, such as rule-based approaches (Weizenbaum, 1966) and statistical models (Brown et al., 1992), generate outputs that resemble human language but do not perfectly reflect the training data. Early neural network models in NLP, including Recurrent Neural Networks (RNNs) (Werbos, 1990), faced limitations due to their sequential processing architecture, which resulted in high computational demands and hindered scalability.
@@
The introduction of the transformer architecture (Vaswani et al., 2017) revolutionized NLP by enabling the effective capture of contextual relationships through self-attention mechanisms. Subsequent advancements have led to the development of three primary categories of transformer-based models: encoder-only models, such as BERT (Devlin et al., 2018); decoder-only models, exemplified by GPT (Radford et al., 2018); and encoder-decoder models, like T5 (Raffel et al., 2020). Decoder-only models, including GPT-4 (Achiam et al., 2023), Claude 3 (Anthropic, 2024), and others (Anil et al., 2023; Team et al., 2023; Touvron et al., 2023), have demonstrated exceptional performance across a wide range of NLP tasks. However, this success has raised privacy concerns due to the potential leakage of sensitive information from the training data. Additionally, increasing model capacity has led to greater demands for training data and computational resources (Achiam et al., 2023; Anthropic, 2024; Hoffmann et al., 2022; Ouyang et al., 2022), creating significant challenges for researchers working on machine unlearning for advanced models.
3. Problem statement
@@
Figure 2. The structure of Iterative Contrastive Unlearning framework. It consists of three parts: (a) Knowledge Unlearning Induction (KUI), (b) Contrastive Learning Enhancement (CLE), and (c) Iterative Unlearning Refinement (IUR).
@@
The architecture of the Iterative Contrastive Unlearning framework, divided into three interconnected modules. Panel (a) (Knowledge Unlearning Induction) unlearns the target data directly. Panel (b) (Contrastive Learning Enhancement) demonstrates paired data processing, where the target data (“Who is Harry Potter?”) undergoes KNN search and the result (“Who is Hermione Granger?”) is used to refine outputs by contrasting correct and incorrect responses, resolving nonsensical repetition seen in intermediate stages. Panel (c) (Iterative Unlearning Refinement) depicts iterative optimization, which adjusts the unlearning process through ongoing evaluation and updates.
4. ICU framework
4.1. Model overview
4.2. Knowledge Unlearning Induction
For a sample in the forget set $\boldsymbol{x}^{\mathrm{fgt}}\in D_{\mathrm{fgt}}$, the sequence of tokens is denoted as $\boldsymbol{x}=(x_{1},x_{2},\ldots,x_{T})$. Following the approach of Jang et al. (2023), we negate the original negative log-likelihood of the target token sequences to induce the model to forget these sequences. The unlearning objective is defined as:
(1) $\mathcal{L}_{\mathrm{fgt}}=\sum_{t=t_{0}}^{T}\log P_{\theta}(x_{t}^{\mathrm{fgt}}|x_{<t}^{\mathrm{fgt}}),$
4.3. Contrastive Learning Enhancement
4.3.1. Analogous data construction.
4.3.2. KNN sampling.
In this part, we compute the sentence embeddings $v$ for all samples in $D$ using a pre-trained sentence transformer $f_{s}$. For each sample $\boldsymbol{x}^{\mathrm{fgt}}$ in $D_{\mathrm{fgt}}$ with its embedding $v_{x}^{\mathrm{fgt}}=f_{s}(\boldsymbol{x}^{\mathrm{fgt}})$, we employ K-Nearest Neighbors (KNN) (Douze et al., 2024) to identify the nearest ($K=1$) embedding $\hat{v_{x}}$ and the corresponding sample $\boldsymbol{\hat{x}}$. All $\boldsymbol{\hat{x}}$ are then collected to form $D_{\mathrm{lrn}}$:
(2) $\boldsymbol{\hat{x}}=\underset{\boldsymbol{x}\in D\setminus D_{\mathrm{fgt}}}{\mathrm{argmin}}\ dis(v_{x},\hat{v_{x}}),$
@@ 4.3. Contrastive Learning Enhancement
4.3.3. Learning enhancement.
The retrieved paired data for $\boldsymbol{x}^{\mathrm{fgt}}$ is denoted as $\boldsymbol{\hat{x}}=\boldsymbol{x}^{\mathrm{lrn}}\in D_{\mathrm{lrn}}$. In contrast to the unlearning objective, we force the model to learn patterns from these paired token sequences using the negative log-likelihood, defined as follows:
(3) $\mathcal{L}_{\mathrm{lrn}}=-\sum_{t=t_{0}}^{T}\log P_{\theta}(x_{t}^{\mathrm{lrn}}|x_{<t}^{\mathrm{lrn}}),$
Additionally, we apply Kullback-Leibler (KL) divergence (Kullback, 1951) to guide the model to approximate the original model’s distribution for data intended to be retained, following Yao et al. (2023):
(4) $\mathcal{L}_{\mathrm{kl}}=\sum_{t=t_{0}}^{T}KL[P_{\theta_{0}}(x_{t}^{\mathrm{lrn}}|x_{<t}^{\mathrm{lrn}})||P_{\theta}(x_{t}^{\mathrm{lrn}}|x_{<t}^{\mathrm{lrn}})],$
@@
4.4. Iterative Unlearning Refinement
@@
(5) $\mathrm{BERTScore}(\boldsymbol{x},f_{\theta}(x_{<t_{0}}))<a,\mathrm{BLEU}(\boldsymbol{x},f_{\theta}(x_{<t_{0}}))<b,$
where $\boldsymbol{x}$ is the referenced target sample, $f_{\theta}(x_{<t_{0}})$ denotes the output of the model provided input sequence $x_{<t_{0}}$, and $a$ and $b$ are predefined thresholds for the iteration process. We empirically determine that a specific token sequence $\boldsymbol{x}$ is considered “forgotten” if Equation (5) is satisfied. Samples deemed “forgotten” are excluded from subsequent epochs. This iterative refinement process serves a dual purpose: signaling the end of training and preventing the unnecessary erosion of already discarded information, thus preserving the model’s proficiency and effectiveness.
Bilingual Evaluation Understudy (BLEU) (Papineni et al., 2002) is a metric originally used to evaluate machine translation quality by measuring the similarity between a model-generated token sequence and one or more reference translations, based on n-gram comparisons. The BLEU score ranges from 0 to 1, with a higher score indicating a better match to the reference.
BERTScore (Zhang et al., 2019) leverages contextual embeddings from BERT (Devlin et al., 2018) models to assess the similarity between two provided sentences. Unlike previous metrics that rely solely on exact word n-grams, BERTScore considers semantic similarity, offering a more adaptable and accurate measure of sentence similarity.
4.5. Training
The objectives in Equation (1), (3) and (4) are jointly used to optimize the model during unlearning. The training process is governed by minimizing the following loss function:
(6) $\mathcal{L}=\mathcal{L}_{\mathrm{fgt}}+\alpha\mathcal{L}_{\mathrm{lrn}}+\beta\mathcal{L}_{\mathrm{kl}},$
@@ 5.1. Datasets
The Pile corpus (825GB) is a large dataset constructed from 22 diverse high-quality subsets, many of which derive from academic or professional sources (e.g. books, open source code) (Gao et al., 2020). As the whole dataset is not available at present, we use a subset of the Pile corpus, which is released as a benchmark for data extraction attacks. [^2] https://github.com/google-research/lm-extraction-benchmark Designed to be easy-to-extract, the subset contains 15,000 samples, randomly sampled from the Pile training dataset. Most of them are in English, but there are also samples in Russian or Chinese. Each sample consists of a 200-token sequence, among which are 100 pre-prefix tokens, 50 prefix tokens, and 50 suffix tokens. Following Jang et al. (2023), we only use the prefix and suffix tokens thus $t_{0}$ is set to 50. This choice of benchmark follows standard practices like Jang et al. (2023) and Kassem et al. (2023), where the same benchmark is widely used to evaluate unlearning methods. While alternative benchmarks such as WMDP (Li et al., 2024) exist, they represent different datasets designed for the same unlearning task, and thus do not fundamentally differ in the nature of the evaluation. Since all these benchmarks focus on generative unlearning tasks, additional experiments on alternative datasets are unnecessary, as they would likely yield similar insights without providing significant new information.
@@
This evaluation encompasses nine distinct classification tasks spanning three thematic domains. Specifically, these domains include linguistic reasoning tasks such as Hellaswag (Zellers et al., 2019) and Lambada (Paperno et al., 2016), as well as assessments of commonsense reasoning through Winogrande (Sakaguchi et al., 2021) and COPA (Gordon et al., 2012). Additionally, scientific reasoning abilities are evaluated through tasks such as ARC-Easy (Clark et al., 2018), ARC-Challenge (Clark et al., 2018), Piqa (Bisk et al., 2020), MathQA (Amini et al., 2019), and PubmedQA (Jin et al., 2019).
Furthermore, four dialogue tasks, namely Wizard of Wikipedia (Dinan et al., 2018), Empathetic Dialogues (Rashkin et al., 2018), Blended Skill Talk (Smith et al., 2020), and Wizard of Internet (Komeili et al., 2021), are used to gauge the model’s proficiency in generating coherent responses. In addition, we measure the perplexity of the unlearned models on the validation set of Pile and Wikitext. Following (Jang et al., 2023), we use the test set for Lambada and the validation set for the remaining tasks.
5.2. Metrics
As stated in Section 5, we assess both learning and unlearning capabilities using two types of datasets. For evaluating unlearning performance, we follow Jang et al. (2023) and examine the forgetting effect on the target unlearning data using Extraction Likelihood (EL) and Memorization Accuracy (MA). In the work of Jang et al. (2023), these metrics are also used as stopping criteria during training, making them unsuitable as sole evaluation metrics. As mentioned in Section 4.3.3, we additionally employ BERTScore and BLEU to measure forgetting during the Iterative Unlearning Refinement (IUR) module.

### v5

# Haoyu Tang1,†, Ye Liu1,†, Xi Zhao2, Xukai Liu1, Yanghai Zhang1 Kai Zhang1,∗ , Xiaofang Zhou2, Enhong Chen1

@@ Abstract
I Introduction
@@
The rapid advancement of deep learning, particularly in Natural Language Processing (NLP), has led to the widespread deployment of powerful generative language models (GLMs) such as GPT-4 [1], Claude 3 [2], and Google Gemini [3]. These models exhibit remarkable capabilities in text generation, summarization, and conversational AI, transforming numerous real-world applications. However, this progress has also raised critical concerns regarding data privacy, security, and compliance with evolving regulations. One of the most pressing challenges is the risk of unintentional memorization and leakage of sensitive information, such as personally identifiable data, proprietary content, and confidential records contained in training datasets [4, 5, 6].
@@
To mitigate these risks, regulatory frameworks like the General Data Protection Regulation (GDPR) [7] have introduced principles such as the “Right To Be Forgotten” (RTBF) [8, 9], which necessitate the ability to remove or selectively forget specific information from trained machine learning models. This has driven extensive research into machine unlearning, a field that aims to ensure models can eliminate particular knowledge while maintaining overall functionality and performance [10, 11].
@@
Figure 1: Difference among the generated sequences of (a) original model, (b) model unlearned by KUMPR [12] and (c) model unlearned by our method.
@@
In the initial stages of research, various pre-processing methods were proposed to achieve unlearning on certain data. For instance, Kandpal et al.[13] found that the likelihood of Generative Language Models regenerating training sequences is correlated with the frequency of those sequences in the training set. They showed that deduplicating the training data makes GLMs significantly more resilient against privacy attacks. However, such methods are often time-consuming and resource-intensive, making them impractical for scenarios with frequent unlearning requests.
@@
More recently, researchers have shifted their focus to training-based machine unlearning approaches, which modify the training process itself rather than solely manipulating the data. For example, SISA [14] partitions the original dataset into several non-overlapping shards and then aggregates models trained on these separate shards. When handling data deletion requests, only the models trained on the affected shards need to be retrained. KGA [15] introduces an additional dataset, using it to fine-tune the original model alongside the original dataset. Nonetheless, both of these methods assume that the training data remain accessible during the unlearning process. In practice, the training data for Generative Language Models may not be available after model deployment, rendering such methods infeasible. A recent method, Knowledge Unlearning for Mitigating Privacy Risks (KUMPR) [12], was designed to address this scenario. [^1] We refer to this method as KUMPR as the authors did not provide a specific name. KUMPR reverses the training objective by maximizing the negative log-likelihood for target tokens and uses metrics to determine whether the model has “forgotten” a target sequence.
@@
  • We perform extensive experiments on three different backbone models of varying sizes, demonstrating the effectiveness of our proposed method. Our code is available at https://github.com/himalalps/ICU.
II Related work
@@ II-A Machine unlearning
Machine unlearning, introduced by Cao et al. [16], aims to protect machine learning models from extraction attacks by removing specific data in such a way that the model behaves as if the data were never part of the training set. Traditional approaches [17] that exclude specific data from training datasets and retrain the model are highly time-consuming and resource-intensive, making them impractical for modern deep neural networks. Similar methods involving retraining [14, 18] also struggle with scalability, particularly when handling numerous deletion requests or when comprehensive datasets are not readily available. To address these challenges, researchers have explored approximate unlearning techniques [19, 20]. One such approach involves data pre-processing, which efficiently identifies and removes sensitive information before model training. Kandpal et al.[13] applied this method to structured private data, such as phone numbers and medical records, and found it effective. However, challenges arise when dealing with unstructured data, as pre-processing may not fully remove all sensitive information [21] and cannot comprehensively address ongoing deletion demands [5].
@@
Recent studies [12, 22, 23, 24, 25] have focused on fine-tuning Generative Language Models to tackle machine unlearning challenges. Jang et al.[12] proposed a novel approach by reversing the traditional training objective, aiming to maximize rather than minimize the negative log-likelihood of tokens designated for forgetting. Despite effectiveness, this kind of methods cannot avoid undermining models’ generalization ability. Other recent methods [26, 27, 28, 22, 15, 29, 23, 30] use diverse techniques such as knowledge gap alignment and reinforcement learning. Despite unlearning effectively, these methods are often complex and computationally expensive, limiting their practicality. For instance, Gu et al.[28] utilized second-order information (Hessian) to provide stronger guarantees for data removal while maintaining model utility, but this approach requires substantial computational resources for Hessian approximation, making it difficult to play a role in real scenarios. Pawelczyk et al.[31] applied in-context methods in unlearning approaches, yet not effective for generation tasks.
Figure 2: The structure of Iterative Contrastive Unlearning framework. It consists of three parts: (a) Knowledge Unlearning Induction (KUI), (b) Contrastive Learning Enhancement (CLE), and (c) Iterative Unlearning Refinement (IUR).
II-B Generative language models
@@
Generative Language Models are designed to understand, generate, and predict human language [32], which have gained considerable attention in recent years.
@@
Traditional language models, such as rule-based approaches [33] and statistical models [34], generate outputs that resemble human language but do not perfectly reflect the training data. Early neural network models in NLP, including Recurrent Neural Networks (RNNs) [35], faced limitations due to their sequential processing architecture, which resulted in high computational demands and hindered scalability.
@@
The introduction of the transformer architecture [36] revolutionized NLP by enabling the effective capture of contextual relationships through self-attention mechanisms [37]. Decoder-only models, including GPT-4 [1], Claude 3 [2], and others [38, 3, 39], have demonstrated exceptional performance across a wide range of NLP tasks [40, 41]. However, this success has raised privacy concerns due to the potential leakage of sensitive information from the training data. Additionally, increasing model capacity has led to greater demands for training data and computational resources [1, 2, 42, 43], creating significant challenges for researchers working on machine unlearning for advanced models.
III Problem statement
@@
IV ICU framework
IV-A Model overview
IV-B Knowledge Unlearning Induction
For a sample in the forget set $\boldsymbol{x}^{\mathrm{fgt}}\in D_{\mathrm{fgt}}$, the sequence of tokens is denoted as $\boldsymbol{x}=(x_{1},x_{2},\ldots,x_{T})$. Following the approach of Jang et al.[12], we negate the original negative log-likelihood of the target token sequences to induce the model to forget these sequences. The unlearning objective is defined as:
$\mathcal{L}_{\mathrm{fgt}}=\sum_{t=t_{0}}^{T}\log P_{\theta}(x_{t}^{\mathrm{fgt}}|x_{<t}^{\mathrm{fgt}}),$ (1)
IV-C Contrastive Learning Enhancement
IV-C1 Analogous data construction.
IV-C2 KNN sampling.
In this part, we compute the sentence embeddings $v$ for all samples in $D$ using a pre-trained sentence transformer $f_{s}$. For each sample $\boldsymbol{x}^{\mathrm{fgt}}$ in $D_{\mathrm{fgt}}$ with its embedding $v_{x}^{\mathrm{fgt}}=f_{s}(\boldsymbol{x}^{\mathrm{fgt}})$, we employ K-Nearest Neighbors (KNN) [44] to identify the nearest ($K=1$) embedding $\hat{v_{x}}$ and the corresponding sample $\boldsymbol{\hat{x}}$. All $\boldsymbol{\hat{x}}$ are then collected to form $D_{\mathrm{lrn}}$:
@@ IV-C Contrastive Learning Enhancement
IV-C3 Learning enhancement.
The retrieved paired data sample for $\boldsymbol{x}^{\mathrm{fgt}}$ is denoted as $\boldsymbol{\hat{x}}=\boldsymbol{x}^{\mathrm{lrn}}\in D_{\mathrm{lrn}}$. In contrast to the unlearning objective, we force the model to learn patterns from these paired token sequences using the negative log-likelihood, defined as follows:
$\mathcal{L}_{\mathrm{lrn}}=-\sum_{t=t_{0}}^{T}\log P_{\theta}(x_{t}^{\mathrm{lrn}}|x_{<t}^{\mathrm{lrn}}),$ (3)
Additionally, we apply Kullback-Leibler (KL) divergence [45] to guide the model to approximate the original model’s distribution for data intended to be retained, following Yao et al.[23]:
$\mathcal{L}_{\mathrm{kl}}=\sum_{t=t_{0}}^{T}KL[P_{\theta_{0}}(x_{t}^{\mathrm{lrn}}|x_{<t}^{\mathrm{lrn}})||P_{\theta}(x_{t}^{\mathrm{lrn}}|x_{<t}^{\mathrm{lrn}})],$ (4)
@@
IV-D Iterative Unlearning Refinement
@@
$\mathrm{BERTScore}(\boldsymbol{x},f_{\theta}(x_{<t_{0}}))<a,$ (5)
$\mathrm{BLEU}(\boldsymbol{x},f_{\theta}(x_{<t_{0}}))<b,$
where $\boldsymbol{x}$ is the referenced target sample, $f_{\theta}(x_{<t_{0}})$ denotes the output of the model provided input sequence $x_{<t_{0}}$, and $a$ and $b$ are predefined thresholds for the iteration process. We empirically determine that a specific token sequence $\boldsymbol{x}$ is considered “forgotten” if (5) is satisfied. Samples deemed “forgotten” are excluded from subsequent epochs. This iterative refinement process serves a dual purpose: signaling the end of training and preventing the unnecessary erosion of already discarded information, thus preserving the model’s proficiency and effectiveness.
Bilingual Evaluation Understudy (BLEU) [46] is a metric originally used to evaluate machine translation quality by measuring the similarity between a model-generated token sequence and one or more reference translations, based on n-gram comparisons. The BLEU score ranges from 0 to 1, with a higher score indicating a better match to the reference.
BERTScore [47] leverages contextual embeddings from BERT [48] models to assess the similarity between two provided sentences. Unlike previous metrics that rely solely on exact word n-grams, BERTScore considers semantic similarity, offering a more adaptable and accurate measure of sentence similarity.
IV-E Training
The objectives in (1), (3) and (4) are jointly used to optimize the model during unlearning. The training process is governed by minimizing the following loss function:
$\mathcal{L}=\mathcal{L}_{\mathrm{fgt}}+\alpha\mathcal{L}_{\mathrm{lrn}}+\beta\mathcal{L}_{\mathrm{kl}},$ (6)
@@ V-A Datasets
The Pile corpus (825GB) is a large dataset constructed from 22 diverse high-quality subsets, many of which derive from academic or professional sources (e.g. books, open source code) [49]. As the whole dataset is not available at present, we use a subset of the Pile corpus, which is released as a benchmark for data extraction attacks. [^2] https://github.com/google-research/lm-extraction-benchmark Designed to be easy-to-extract, the subset contains 15,000 samples, randomly sampled from the Pile training dataset. Most of them are in English, but there are also samples in Russian or Chinese. Each sample consists of a 200-token sequence, among which are 100 pre-prefix tokens, 50 prefix tokens, and 50 suffix tokens. Following Jang et al.[12], we only use the prefix and suffix tokens thus $t_{0}$ is set to 50. This choice of benchmark follows standard practices like Jang et al.[12] and Kassem et al.[22], where the same benchmark is widely used to evaluate unlearning methods. While alternative benchmarks such as WMDP [30] exist, they represent different datasets designed for the same unlearning task, and thus do not fundamentally differ in the nature of the evaluation. Since all these benchmarks focus on generative unlearning tasks, additional experiments on alternative datasets are unnecessary, as they would likely yield similar insights without providing significant new information.
@@
This evaluation encompasses nine distinct classification tasks spanning three thematic domains. Specifically, these domains include linguistic reasoning tasks such as Hellaswag [50] and Lambada [51], as well as assessments of commonsense reasoning through Winogrande [52] and COPA [53]. Additionally, scientific reasoning abilities are evaluated through tasks such as ARC-Easy [54], ARC-Challenge [54], Piqa [55], MathQA [56], and PubmedQA [57].
Furthermore, four dialogue tasks, namely Wizard of Wikipedia [58], Empathetic Dialogues [59], Blended Skill Talk [60], and Wizard of Internet [61], are used to gauge the model’s proficiency in generating coherent responses. In addition, we measure the perplexity of the unlearned models on the validation set of Pile and Wikitext. Following [12], we use the test set for Lambada and the validation set for the remaining tasks.
V-B Metrics
As stated in Section V, we assess both learning and unlearning capabilities using two types of datasets. For evaluating unlearning performance, we follow Jang et al.[12] and examine the forgetting effect on the target unlearning data using Extraction Likelihood (EL) and Memorization Accuracy (MA). In the work of Jang et al.[12], these metrics are also used as stopping criteria during training, making them unsuitable as sole evaluation metrics. As mentioned in Section IV-C3, we additionally employ BERTScore and BLEU to measure forgetting during the Iterative Unlearning Refinement (IUR) module.